OpenStudy (anonymous):

Can someone please help! Find the focus of the parabola that has vertex (-4,-1), opens vertically, and passes through the point (-3,4).

OpenStudy (dumbcow):

ok for a given parabola \[y = a(x-h)^{2} +k\] the focus is \[(h , k+ \frac{1}{4a})\] you are given the vertex (h,k) , you can solve for "a" using point (-3,4)

OpenStudy (anonymous):

just unsure of how to work (-3,4) into the problem.. hmm

OpenStudy (anonymous):

(-4,-15/16) ?

OpenStudy (anonymous):

nope it must be (-4,-19/20)... but why did I get -15/20? hmm

OpenStudy (anonymous):

-15/16 i mean

OpenStudy (dumbcow):

oh you substitute (-3,4) as (x and y in parabola equation --> 4 = a(-3+4)^2 -1 solve for a

OpenStudy (anonymous):

right and i get 4

OpenStudy (dumbcow):

hmm 4 = a(1)^2 -1 4 = a-1 5 = a

OpenStudy (anonymous):

ohh i was getting 1^2-1=0 ... but i get it now

OpenStudy (anonymous):

so -19/20 got it! thanks!

OpenStudy (dumbcow):

yw